Solve by addition or substitution.
3x + 5y = 2
2x + 5y = -2

Answer by elima(1433)

We will solve this by addition;
3x + 5y = 2
(-1)2x + 5y = -2
====================
3x +5y = 2
-2x -5y = 2
------------
x=4
Now that we have our x lets plug that in to find y;
3(4)+5y=2
12 +5y =2
5y = -10
y = -2
=============
:)